# Spiking Transformers Need High Frequency Information

### v1

@@ 5 Conclusion
In Table 6, we provide additional analysis on the impact of high-frequency information. We conducted these experiments on CIFAR-100 [35], using the experimental settings detailed in Table 5. We discuss the ablation below according to the following aspects:
(a) Extra Max-Pooling in Patch Embedding/Token Mixing:
@@
Appendix A2 Energy Analysis
To estimate the theoretical energy consumption of Spiking Transformers, we follow the methodology used in previous studies [7, 49, 39, 18]. It is worth noting that the batch normalization (BN) layers and linear scaling transformations that follow convolution layers can be combined directly into the convolution layers themselves with added bias terms during deployment. Thus, in common practice [7, 49, 39, 18], the energy consumption of BN is typically excluded when calculating theoretical energy usage. For fair comparison, our work adopts the same strategy. In Spiking Transformers, energy consumption is directly proportional to synaptic operations (SOPs), which can be calculated as:
@@
where $l$ represents a specific block or layer in the Spiking Transformer architecture, $fr$ refers to the firing rate of the input spike train for that particular block or layer, and T is the simulation time step of spiking neurons. Assuming the multiply-accumulate (MAC) and accumulate (AC) operations are implemented on the 45nm neuromorphic chip described in [6], where each MAC operation uses $E_{\text{MAC}}$ = 4.6pJ of energy and each AC operation uses $E_{\text{AC}}$ = 0.9pJ, we can estimate the total energy consumption of a Spiking Transformer by adding up the energy used by all MAC and AC operations across all layers:
@@
Appendix A3 Comparison on Train/Inference Time and Memory Consumption
Table 7: Comparison of training/inference time and memory usage between QKFormer-10-768 and Max-Former-10-768 models. All measurements were conducted with the simulation timestep of 1 and the batch size of 32. QKFormer* indicates the QKFormer variant with membrane shortcut.
@@
QKFormer* (64.96M, T=1, B=32) 0.208 17496 0.048 4822
Max-Former (63.99M, T=1, B=32) 0.179 15431 0.044 4354
Max-Former delivers faster training and inference speeds while consuming less memory. We compared the performance metrics of QKFormer [22], its membrane shortcut variant (QKFormer*), and Max-Former on ImageNet using 224 $\times$ 224 input resolution. All tests were conducted on a CentOS 7.9 server equipped with the Intel Xeon Gold 6348 CPU (2.60GHz) and the Nvidia A30 GPU. As Table 7 shows, when compared to QKFormer* with the same hierarchical architecture and shortcut configuration, Max-Former reduces training time by 14% and both inference time and memory usage by 10%. Additionally, our results indicate that the pre-spike shortcut strategy used in the original QKFormer increases both processing time and memory demands.
Appendix A4 Residual Connections in Spiking Neural Networks
@@
The unique asynchronous nature of spike-based computation makes implementing residual connections challenging. As a result, the research community of spiking neural networks (SNNs) has not yet reached a consensus on the standardized residual learning approach, either in terms of algorithm or hardware implementation. While the main focus of our work is not related to residual learning, we want to offer a detailed comparison between the pre-spike shortcut [50] and the membrane shortcut [49], the two most representative residual learning methods that emerged in recent years.
@@
As explained in Section 3.3, the pre-spike shortcut [50] implements residual connections between spiking outputs, while the membrane shortcut [49] connects membrane potentials directly. In algorithm designs, membrane shortcuts have been reported to facilitate better performance in prior works [22, 39], especially on small datasets. However, our findings indicate this advantage is not universal across all scenarios. As shown in Table 8, the patch embedding stages of Spiking Transformers account for the majority of energy consumption. Consequently, the multiple patch embedding stages in hierarchical architectures like QKFormer  [22], while enabling efficient feature learning with fewer parameters, also come at higher energy usage. This makes the choice of shortcut scheme particularly impactful on overall energy efficiency. When processing ImageNet images at 224 × 224 resolution, the pre-spike QKFormer consumes three times more energy than its membrane shortcut variant, reflecting the substantially higher SOPs required. Nevertheless, this increased computational overhead does translate to notable performance gains (+2.32% accuracy when comparing QKFormer to QKFormer*).
@@
From the hardware perspective, implementing either shortcut type on neuromorphic chips is technically feasible but presents significant challenges [51, 52]. Implementing the pre-spike shortcut specifically requires the chips to support multi-spike operations, as the ternary spike transmissions (values of 0, 1, or 2) can occur. This results in either higher energy consumption or increased hardware complexity [53]. Yao et al. [39] proposed implementing membrane shortcuts through the addressing function that passes the membrane potential to corresponding neurons in subsequent layers for merging. While membrane shortcuts do strictly adhere to the spike-driven computing paradigm and could possibly be supported by standard neuromorphic hardware, transmitting membrane potentials can create substantial communication overhead, making practical implementation non-trivial. Many discussions on hardware deployment still advocate avoiding shortcuts as the current preferred approach [54, 55]. However, given the critical role of residual learning in modern deep learning, avoiding shortcuts altogether is not a sustainable long-term strategy for advancing neuromorphic computing.
@@
In our work, we primarily compare with QKFormer* using the membrane shortcut to ensure fair comparisons. We hope the SNN community can establish a consensus on standardized shortcut implementations in the near future, considering the significant impact of shortcut schemes on both energy efficiency and model performance.
@@
Table 8: Energy consumption comparison (in mJ) across processing stages for QKFormer, QKFormer with membrane shortcut (QKFormer*), and our proposed Max-Former model.
@@
QKFormer* (16.47M, T=4, 76.48%) 1.19 0.052 0.15 0.96 0.11 0.25 0.88 0.88 1.06 0.007 5.52
Max-Former (16.23M, T=4, 77.82%) 0.41 0.02 0.17 0.89 0.01 0.36 0.91 0.96 1.16 0.001 4.89
(b) Visualization:
@@
We demonstrate the GradCAM visualizations [56] of four Spiking Transformers with the membrane shortcut of similar size. Compared to Spike-Driven Transformer [39] and SWFormer [18], the hierarchical architecture used in both QKFormer [22] and our Max-Former allows for more precise focusing on target objects. Compared to QKFormer*, Max-Former shows more concentrated activation patterns. For instance, in the polar bear image, Max-Former completely skips the background and precisely focuses on the bear’s key features (the head, rather than the outline or fur).
@@
Figure 7: GradCAM visualizations [56] comparing four Spiking Transformers of similar size: Spike-Driven Transformer-8-512 [39] (29.68M), SWFormer-8-512 [18] (27.6M), QKFormer-10-512 with membrane shortcut (QKFormer*) [22] (29.08M), and our Max-Former-10-512 (28.65M).
Appendix A5 Limitation
@@
In this work, we provide theoretical evidence that high-frequency information is essential for Spiking Transformers. To empirically probe the importance of high-frequency information, we design Max-Former. However, Max-Former should be considered just a good starting baseline for future architectural designs. Similar to [29], Max-Former requires manually balancing frequency components, which demands considerable expertise when adapting to different tasks. Incorporating direct frequency learning approaches like Fourier-based [57] or Wavelet-based[18] methods would offer more straightforward solutions. The main challenge, however, lies in developing efficient spike-based frequency representations. Overall, we believe our work will inspire more future research to advance neuromorphic computing through exploring the unique properties of spiking neurons, rather than expending excessive effort to adapt established practices from standard non-spiking neural networks.

### v6

@@ A.2 Impact of High-Frequency Information on Model Performance
(6) Embed-Orig/Embed-Max/Embed-Max SSA/SSA/SSA 82.13
Embed-1Max/Embed-Max/Embed-Max SSA/SSA/SSA 79.98
Embed-Max/Embed-Max/Embed-Max SSA/SSA/SSA 78.78
(7) Embed-Orig/Embed-1Max/Embed-1Max SSA/SSA/SSA 82.02
Embed-1Max/Embed-1Max/Embed-1Max SSA/SSA/SSA 81.86
(8) Embed-Orig/Embed-Max/Embed-Max SDSA/SDSA/SDSA 81.77
Embed-Orig/Embed-1Max/Embed-1Max SDSA/SDSA/SDSA 81.52
In Table 7, we provide additional analysis on the impact of high-frequency information. We conducted these experiments on CIFAR-100 [krizhevskyLearningMultipleLayers2009a], using the experimental settings detailed in Table 6. We discuss the ablation below according to the following aspects:
(a) Extra Max-Pool in Patch Embedding/Token Mixing:
@@
In general, for the hierarchical spiking transformer, Embed-Max can improve performance when used for patch embedding, as shown in Table 7 (6-8). However, for non-hierarchical ones, although enhancing high-frequency components can in principle be beneficial, its actual impact is limited since patch embedding is only performed once throughout the entire process, causing the improvement to diminish in deeper layers. A more promising solution lies in optimizing token mixing, as discussed in [fang2024spiking].
@@
A.3 Max-ResNet Implementation
As shown in Figure 7, Max‑ResNet introduces only a minor architectural change to MS‑ResNet [hu2021advancing]: all layers are replaced with Max‑ResNet layers, while the first layer remains unchanged. Code implementation is available at https://github.com/bic-L/MaxFormer.
Figure 7: Overview of Max-ResNet. A single Max‑Pool operation is added per block and layer.
Appendix B Energy Analysis
To estimate the theoretical energy consumption of Spiking Transformers, we follow the methodology used in previous studies [zhou2022spikformer, hu2021spiking, yao2023spike, fang2024spiking]. It is worth noting that the batch normalization (BN) layers and linear scaling transformations that follow convolution layers can be combined directly into the convolution layers themselves with added bias terms during deployment. Thus, in common practice [zhou2022spikformer, hu2021spiking, yao2023spike, fang2024spiking], the energy consumption of BN is typically excluded when calculating theoretical energy usage. For fair comparison, our work adopts the same strategy. In Spiking Transformers, energy consumption is directly proportional to synaptic operations (SOPs), which can be calculated as:
@@ Appendix B Energy Analysis
where $l$ represents a specific block or layer in the Spiking Transformer architecture, $fr$ refers to the firing rate of the input spike train for that particular block or layer, and T is the simulation time step of spiking neurons. Assuming the multiply-accumulate (MAC) and accumulate (AC) operations are implemented on the 45nm neuromorphic chip described in [davies2018loihi], where each MAC operation uses $E_{\text{MAC}}$ = 4.6pJ of energy and each AC operation uses $E_{\text{AC}}$ = 0.9pJ, we can estimate the total energy consumption of a Spiking Transformer by adding up the energy used by all MAC and AC operations across all layers:
@@
Appendix C Comparison on Train/Inference Time and Memory Consumption
Table 8: Comparison of training/inference time and memory usage between QKFormer-10-768 and Max-Former-10-768 models. All measurements were conducted with the simulation timestep of 1 and the batch size of 32. MS-QKFormer indicates the QKFormer variant with the membrane shortcut.
@@ Appendix C Comparison on Train/Inference Time and Memory Consumption
MS-QKFormer* (64.96M, T=1, B=32) 0.208 17496 0.048 4822
Max-Former* (63.99M, T=1, B=32) 0.179 15431 0.044 4354
Max-Former delivers faster training and inference speeds while consuming less memory. We compared the performance metrics of QKFormer [zhou2024qkformer], its membrane shortcut variant (MS-QKFormer), and Max-Former on ImageNet using 224 $\times$ 224 input resolution. All tests were conducted on a CentOS 7.9 server equipped with the Intel Xeon Gold 6348 CPU (2.60GHz) and the Nvidia A30 GPU. As Table 8 shows, when compared to MS-QKFormer with the same hierarchical architecture and shortcut configuration, Max-Former reduces training time by 14% and both inference time and memory usage by 10%. Additionally, our results indicate that the pre-spike shortcut strategy used in the original QKFormer increases both processing time and memory demands.
Appendix D Residual Connections in Spiking Neural Networks
@@
The unique asynchronous nature of spike-based computation makes implementing residual connections challenging. As a result, the research community of spiking neural networks (SNNs) has not yet reached a consensus on the standardized residual learning approach, either in terms of algorithm or hardware implementation. While the main focus of our work is not related to residual learning, we want to offer a detailed comparison between the pre-spike shortcut [fang2021deep] and the membrane shortcut [hu2021spiking], the two most representative residual learning methods that emerged in recent years.
@@
As explained in Section 3.3, the pre-spike shortcut [fang2021deep] implements residual connections between spiking outputs, while the membrane shortcut [hu2021spiking] connects membrane potentials directly. In algorithm designs, membrane shortcuts have been reported to facilitate better performance in prior works [zhou2024qkformer, yao2023spike], especially on small datasets. However, our findings indicate this advantage is not universal across all scenarios. As shown in Table 9, the patch embedding stages of Spiking Transformers account for the majority of energy consumption. Consequently, the multiple patch embedding stages in hierarchical architectures like QKFormer  [zhou2024qkformer], while enabling efficient feature learning with fewer parameters, also come at higher energy usage. This makes the choice of shortcut scheme particularly impactful on overall energy efficiency. When processing ImageNet images at 224 × 224 resolution, the pre-spike QKFormer consumes three times more energy than its membrane shortcut variant, reflecting the substantially higher SOPs required. Nevertheless, this increased computational overhead does translate to notable performance gains (+2.32% accuracy when comparing QKFormer to MS-QKFormer).
@@
From the hardware perspective, implementing either shortcut type on neuromorphic chips is technically feasible but presents significant challenges [weidel2021wavesense, balafrej2025enhancing]. Implementing the pre-spike shortcut specifically requires the chips to support multi-spike operations, as the ternary spike transmissions (values of 0, 1, or 2) can occur. This results in either higher energy consumption or increased hardware complexity [gautam2023adaptive]. Yao et al. [yao2023spike] proposed implementing membrane shortcuts through the addressing function that passes the membrane potential to corresponding neurons in subsequent layers for merging. While membrane shortcuts do strictly adhere to the spike-driven computing paradigm and could possibly be supported by standard neuromorphic hardware, transmitting membrane potentials can create substantial communication overhead, making practical implementation non-trivial. Many discussions on hardware deployment still advocate avoiding shortcuts as the current preferred approach [deng2024spiking, meyer2024diagonal]. However, given the critical role of residual learning in modern deep learning, avoiding shortcuts altogether is not a sustainable long-term strategy for advancing neuromorphic computing.
@@
In our work, we primarily compare with MS-QKFormer using the membrane shortcut to ensure fair comparisons. We hope the SNN community can establish a consensus on standardized shortcut implementations in the near future, considering the significant impact of shortcut schemes on both energy efficiency and model performance.
@@
Table 9: Energy consumption comparison (in mJ) across processing stages for QKFormer, QKFormer with membrane shortcut (MS-QKFormer), and our proposed Max-Former model. * Identical training configurations.
@@
MS-QKFormer* (16.47M, T=4, 76.48%) 1.19 0.052 0.15 0.96 0.11 0.25 0.88 0.88 1.06 0.007 5.52
Max-Former* (16.23M, T=4, 77.82%) 0.41 0.02 0.17 0.89 0.01 0.36 0.91 0.96 1.16 0.001 4.89
Appendix E Visualization
@@
We demonstrate the GradCAM visualizations [selvarajuGradcamVisualExplanations2017] of four Spiking Transformers with the membrane shortcut of similar size. Compared to Spike-Driven Transformer [yao2023spike] and SWFormer [fang2024spiking], the hierarchical architecture used in both QKFormer [zhou2024qkformer] and our Max-Former allows for more precise focusing on target objects. Compared to MS-QKFormer, Max-Former shows more concentrated activation patterns. For instance, in the polar bear image, Max-Former completely skips the background and precisely focuses on the bear’s key features (the head, rather than the outline or fur).
@@
Figure 8: GradCAM visualizations [selvarajuGradcamVisualExplanations2017] comparing four Spiking Transformers of similar size: Spike-Driven Transformer-8-512 [yao2023spike] (29.68M), SWFormer-8-512 [fang2024spiking] (27.6M), QKFormer-10-512 with membrane shortcut (MS-QKFormer) (29.08M), and our Max-Former-10-512 (28.65M).
Appendix F Impact and Limitation
Our work serves as the theoretical foundation for many existing architectural choices in Spiking Transformers. Specifically, in  [yao2023spike], the authors found that directly applying known practices of MetaFormer does not achieve good results in Spiking Transformers: employing average pooling operators to replace SDSA as the token mixer surprisingly results in substantial performance degradation from 61.0% to 41.2%. Similar phenomena have been discussed in earlier works. For instance, Spikformer v2 [zhou2024spikformer] discovered that removing the max-pooling operator in the original Spikformer [zhou2022spikformer] leads to a substantial performance drop, while adding convolution layers (which act as high-pass filters) in the patch embedding stage significantly improves the performance. Our work reveals the underlying principles of these architectural designs: spiking transformers need to enhance high-frequency components to alleviate the feature degradation caused by their inherent low-pass activation.
@@ Appendix F Impact and Limitation
We are well aware that there is still much space to be explored, and we hope that our Max-Former can serve as a good starting point for future research. For instance, similar to [si2022inception], Max-Former requires manually balancing frequency components, which demands considerable expertise when adapting to different tasks. Incorporating direct frequency learning approaches like Fourier-based [li2020fourier] or Wavelet-based[fang2024spiking] methods would offer more straightforward solutions. The main challenge, however, lies in developing efficient spike-based frequency representations. Overall, we believe our work will inspire more future research to advance neuromorphic computing through exploring the unique properties of spiking neurons, rather than expending excessive effort to adapt established practices from standard non-spiking neural networks.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction reflect the paper’s contributions and scope clearly.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We discuss the limitations in Appendix E.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: All the theorems, formulas, and proofs are clearly stated in the main paper.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We provide experiment details in the Appendix AB.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. 6.
    Experimental setting/details
  21. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  22. Answer: [Yes]
  23. Justification: All training and test details have been revealed in the Appendix A.
  24. 7.
    Experiment statistical significance
  25. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  26. Answer: [No]
  27. Justification: Error bars are not reported because it would be too computationally expensive.
  28. 8.
    Experiments compute resources
  29. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  30. Answer: [Yes]
  31. Justification: See Appendix A.
  32. 9.
    Code of ethics
  33. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  34. Answer: [Yes]
  35. Justification: This research is in every respect with the NeurIPS Code of Ethics.
  36. 10.
    Broader impacts
  37. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  38. Answer: [N/A]
  39. Justification: This research belongs to foundational research and is not tied to particular applications.
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  42. Answer: [N/A]
  43. Justification: The paper poses no such risks.
  44. 12.
    Licenses for existing assets
  45. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  46. Answer: [Yes]
  47. Justification: The paper properly credits creators and mentions the license and terms of use for existing assets.
  48. 13.
    New assets
  49. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  50. Answer: [N/A]
  51. Justification: The paper does not release new assets.
  52. 14.
    Crowdsourcing and research with human subjects
  53. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  54. Answer: [N/A]
  55. Justification: The paper does not involve crowdsourcing or research with human subjects.
  56. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  57. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  58. Answer: [N/A]
  59. Justification: The paper does not involve crowdsourcing or research with human subjects.
  60. 16.
    Declaration of LLM usage
  61. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  62. Answer: [N/A]
  63. Justification: The paper does not involve LLMs as any important, original, or non-standard components.